Superfluid Density and Condensation of Exotic Superconductors: Bose-Einstein, Kosterlitz-Thouless, BCS, and else ?

Y.J. Uemura

Physics Department, Columbia University, New York, NY 10027, USA

Since 1987, we have measured the muon spin relaxation rate σ in high-T_{c} cuprates and other exotic superconductors to determine the magnetic field penetration depth λ. We noticed universal correlations between T_{c} and σ(T → 0) ∝ 1/λ^{2} ∝ n_{s}/m^{∗} (superconducting carrier density / effective mass) in underdoped cuprates [1]. Combining this observation with the “pseudo-gap” behavior, we proposed a picture for high-T_{c} cuprates based on crossover from Bose-Einstein to BCS condensation [2,3].

In this talk, we would like to consider relationship of these results and picture to Kosterlitz-Thouless transition. For this purpose, we examine (1) models regarding phase fluctuations [4]; (2) variation of T_{c} in YBCO-PrBCO films [5] and Bi2212 system intercalated with organic molecules [6]; and (3) recent high-frequency conductivity studies on Bi2212 [7]:

We will also discuss similarities of HTSC, organic BEDT, A_{3}C_{60} and intercalated HfNCl superconductors. In particular, we notice that superfluid density n_{s}/m^{∗} diminishes when the system crosses over from superconductor to simpler metal by overdoping / application of pressure, contrary to the behavior expected for BCS superconductors.

[1] Y.J. Uemura et al., Phys. Rev. Lett. 62 (1989) 2317; Phys. Rev. Lett. 66 (1991) 2665.

[2] Y.J. Uemura, in Proceedings of the Workshop in Polarons and Bipolarons in High-T_{c} Superconductors and Related Materials, Cambridge, 1994, ed. by E. Salje et al. (Cambridge Univ. Press, 1995), p.p. 453.

[3] Y.J. Uemura, Physica C 282-287 (1997) 194.

[4] V.J. Emery and S. Kivelson, Nature 374 (1995) 434.

[5] O. Fisher et al., Physica B169 (1991) 116.

[6] J.H. Choy et al., Science 280 (1995) 1589.

[7] J. Corson et al., Nature 389 (1999) 221.